Given the following base compositions, determine which double-stranded DNA molecule is expected to be the most thermodynamically stable?

■A + T = 10%

■A + T = 30%

■G + C = 60%

■G + C = 50%

G + C = 40%

Please explain!!
Answer #1

Thermostability of DNA :-
• When temperature is increased hydrogen bonds between nitrogen bases of complementary DNA strands broken.
• DNA contains four bases- Guanine, Cytosine, Adenine, and Thymine. Hydrogen bonds found between these bases.
• Double hydrogen bonds present between adenine and thymine while triple hydrogen bonds present between cytosine and guanine.
• If there is more number of hydrogen bonds present in DNA, then more heat will require for bond degradation and this DNA will be more thermostable.

Now, which DNA has more hydrogen bonds.
• A DNA with more number of cytosine and guanine has more hydrogen bonds (*because triple hydrogen bonds present between them which is more then double bonds between adenine and thymine).
• So, A DNA strand with more Cytosine and guanine has higher thermostability.
More G-C amount in DNA = More thermostable.

Now in question,
(A+T) + (G+C) = 100%
GC amount in DNAs -
(a) A+T = 10% so G+C = 100-10 = 90%
(b) A+T = 30% so G+C = 70%.
(c) G+C = 60%.
(d) G+C = 50%
(e) G+C = 40%.

Option (a) is most thermostable because highest amount of G+C.
Thermostability order -
a>b>c>d>e